OpenStudy (anonymous):

Find the equation for the inverse of each equation: f(x)=x^2-4

OpenStudy (anonymous):

I have this as an answer X+4=F^-1(X)

OpenStudy (zehanz):

Write y=x²-4, let x and y switch position: x=y²-4 and solve for y...

OpenStudy (zehanz):

Your answer is wrong I'm afraid... If x=y²-4 then y²=x+4, so y=...?

OpenStudy (anonymous):

I have this so far \[\sqrt{X+4} = \sqrt{Y^2}\]

OpenStudy (zehanz):

Yes, so\[y=\sqrt{x+4}\]or better:\[f^{-1}(x)=\sqrt{x+4}\]

OpenStudy (anonymous):

So then that would be the answer after i i have squared both sides right? \[F^-1 =\sqrt{X+4}\]

OpenStudy (zehanz):

See my last answer above... BTW, you really should avoid using the capitals...

OpenStudy (zehanz):

They don't even fit under the radicals ;)

OpenStudy (anonymous):

Thanks:D & okay i will^-^

OpenStudy (zehanz):

YW!